A fairgrounds ride spins its occupants inside a flying-saucer-shaped container. If the horizontal circular path the riders follow has a 6.00 m radius, at how many revolutions per minute will the riders be subjected to a centripetal acceleration 2.95 times that of gravity?
Answer #1

Acceleration,a

a = 2.95*9.8 = 28.91 m/s^2

a = v^2/r = w^2*r

w = sqrt(a/r)

= sqrt(28.91/6)

= 2.2 rad/s

= 2.2*60/(2*pi) rev/min

= 20.96 rev/min